OpenStudy (anonymous):

Calculate the derivative: dP/dV where v=-2 and p=4/v no idea how to start this one.

hartnn (hartnn):

can u calculate d/dv(v^n) ?

OpenStudy (anonymous):

\[nx ^{n-1}\]

hartnn (hartnn):

right :) \(so\quad d/dv(1/v) = d/dv(v^{-1})=?\)

OpenStudy (anonymous):

-v^-2?

hartnn (hartnn):

that is correct ! :) so \(d/dv(4/v)= 4 d/dv(1/v)=?\)

hartnn (hartnn):

because 4 is a constant and can be taken out from derivative

OpenStudy (anonymous):

=4

hartnn (hartnn):

dP/dv = d/dv(4/v) = 4 d/dv(1/v) = 4 \((-v^{-2})\) got this ? now just put v= -2 here

OpenStudy (anonymous):

??

OpenStudy (anonymous):

that all equals one, right?

OpenStudy (anonymous):

4(-v^-2) = 4(2^-2)=4 times 1/4 = 1

hartnn (hartnn):

\(\huge4(-(-2)^{(-2)})= -4(\frac{1}{(-2)^2})=\frac{-4}{4}=?\) u just missed a - *minus* sign

OpenStudy (anonymous):

shouldn't that be a pus 2?

OpenStudy (anonymous):

-v = -(-2) = 2

hartnn (hartnn):

but the derivative, does not contain -v derivative is \(\huge -v^{-2}= -\frac{1}{v^2}\)

OpenStudy (anonymous):

so the derivative of v=-1

hartnn (hartnn):

derivative of v will be +1 but why do u need derivative of v ?

OpenStudy (anonymous):

the question is the derivative of dP/dV and you told me to get the derivative of v .. and i said it was plus one then you said i was wrong and that it's -4/4

OpenStudy (anonymous):

im just trying to find out the answer haha.

hartnn (hartnn):

i never said that derivative of v , +1 is wrong....

OpenStudy (anonymous):

im confused as to what you are trying to get me to do then.

hartnn (hartnn):

i just said that u missed a minus sign here "4(-v^-2) = 4(2^-2)=4 times 1/4 = 1 " it should be 4(-v^-2) = 4(-(-2)^-2)=4 times (-1/4) = -1 got this ?

hartnn (hartnn):

sorry, if i confused you....

OpenStudy (anonymous):

and is that the derivative for v?

hartnn (hartnn):

u need to calculate derivative of P, dP/dv which we worked out as -1 and thats the final answer we don't need derivative of v here(which is 1 bdw)

OpenStudy (anonymous):

so my answer is -1?

hartnn (hartnn):

YES ! :)

OpenStudy (anonymous):

thank you so much !!!

OpenStudy (anonymous):

thankss

hartnn (hartnn):

welcome ^_^